OpenStudy (anonymous):

i dont get it!!! im so confused...HELP!!!:( FOIL method (x-2)(x+3)

OpenStudy (pottersheep):

This video will help you tons! http://www.youtube.com/watch?v=65BAsV0SK3k

jimthompson5910 (jim_thompson5910):

FOIL stands for: First, Outer, Inner, Last First: first terms are x and x, they multiply to x*x = x^2 Outer: outer terms are x and 3, they multiply to x*3 = 3x Inner: inner terms are -2 and x, they multiply to -2*x = -2x Last: last terms are -2 and 3, they multiply to -2*3 = -6 Now add up the last 4 results to get x^2 + 3x + (-2x) + (-6) x^2 + 3x - 2x - 6 x^2 + 1x - 6 x^2 + x - 6

jimthompson5910 (jim_thompson5910):

So (x-2)(x+3) expands out to x^2 + x - 6